YoursTruly:

Solve by Completing the Square. 2x^2 + 9x + 4 = 0 x=-2 and x=-4 x=2 and x=9 x=-1/2 and x=-4

YoursTruly:

@Allison @Ultrilliam @JustSaiyan

psirockin2:

2x^2 -9x + 4 = 0 x^2 -(9/2)x = -2 x^2 -(9/2)x + (9/4)^2 = -2 + (9/4)^2 (x-(9/4))^2 = -32/16 + 81/16 (x-(9/4))^2 = 49/16 x-(9/4) = +-7/4 x = +(7/4)+(9/4) or x = -(7/4)+(9/4) x = *-1/2* 4 x^2 + x = 0 x^2 + (1/4)x = 0 x^2 + (1/4)x + (1/8)^2 = (1/8)^2 (x+(1/8))^2 = (1/8)^2 (x + (1/8)) = +-(1/8) x = (1/8)-(1/8) or x = -(1/8)-(1/8) x = *-1/4* So I'm pretty sure it's C

YoursTruly:

Thank youuuuu <3

psirockin2:

np